# Existence of cohomogeneity one Einstein metrics

### v4

@@ 5. Appendix
(5.1) $\begin{split}-d_{1}^{2}(n-1)(P_{X}+Q_{X})&=(d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}-d_{2}^{3}+d_{1}d_{2}+d_{2}^{2})k^{3}\\ &\quad+(d_{1}^{2}d_{2}^{2}+2d_{1}^{2}d_{2}-d_{1}d_{2}^{2}+d_{1}d_{2})k^{2}\\ &\quad+(d_{1}^{3}d_{2}-d_{1}^{2}d_{2})k+(2d_{1}^{3}-2d_{1}^{2})(k-k^{2})+d_{1}^{3}(d_{1}-1)(1-k).\end{split}$
The last four terms above are non-negative and do not vanish simultaneously for any $k\in[0,1]$. If the coefficient of $k^{3}$ is non-negative, then the polynomial is positive on $[0,1]$. If the coefficient of $k^{3}$ is negative, we have
(5.2) $\begin{split}-d_{1}^{2}(n-1)(P_{X}+Q_{X})&>(d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}-d_{2}^{3}+d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(d_{1}^{2}d_{2}^{2}+2d_{1}^{2}d_{2}-d_{1}d_{2}^{2}+d_{1}d_{2})k^{2}\\ &=2d_{1}d_{2}(d_{1}d_{2}-d_{1}-d_{2})k^{2}+(d_{2}^{2}-d_{2})(d_{1}d_{2}-d_{1}-d_{2})k^{2}+d_{1}d_{2}k^{2}.\\ \end{split}$
Given $d_{2}\geq d_{1}\geq 2$, the polynomial above is non-negative on $[0,1]$. Hence, we have $P_{X}+Q_{X}<0$ on $[0,1]$. ∎
Proposition 5.2.
For any $d_{2}\geq d_{1}\geq 2$, the inequalities $\beta_{0}(k)>0$ and $\beta_{1}(k)>0$ hold for any $k\in[0,1]$.
@@
(5.4) $\begin{split}\beta_{1}(k)&=(2d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+2d_{2})k^{2}\\ &\quad+(4d_{1}^{3}d_{2}+2d_{1}^{2}d_{2}^{2}-4d_{1}^{3}-2d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+4d_{1}d_{2}+d_{2}^{2}+4d_{1}-2d_{2})k\\ &\quad+2d_{1}^{4}+d_{1}^{3}d_{2}-2d_{1}^{2}d_{2}+2d_{1}^{2}+d_{1}d_{2}-4d_{1}.\\ \end{split}$
If $d_{1}=d_{2}=2$, then $\beta_{1}(k)=-12k^{2}+48k+36$ is positive on $[0,1]$. For $d_{2}\geq 3$, we argue that the coefficient of each $k^{i}$ is positive. The constant term of $\beta_{1}(k)$ is apparently positive. The coefficient of $k$ is
(5.5) $\begin{split}&4d_{1}^{3}d_{2}+2d_{1}^{2}d_{2}^{2}-4d_{1}^{3}-2d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+4d_{1}d_{2}+d_{2}^{2}+4d_{1}-2d_{2}\\ &=2d_{1}^{2}(2d_{1}d_{2}-2d_{1}-d_{2})+d_{1}d_{2}(2d_{1}d_{2}-3d_{2})+4d_{1}d_{2}+(d_{2}^{2}-2d_{2})+4d_{1},\end{split}$
which is positive for any $d_{2}\geq d_{1}\geq 2$. The coefficient of $k^{2}$ is
(5.6) $\begin{split}&2d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-4d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+2d_{2}\\ &=(2d_{2}^{2}-4d_{2})d_{1}^{2}+(d_{2}^{3}-2d_{2}^{2}-2d_{2})d_{1}-d_{2}^{3}+2d_{2}.\end{split}$
As a function of $d_{1}$, the expression above is positive at $d_{1}=2$ and increasing on $[2,\infty)$. Therefore, we have $\beta_{1}>0$ on $[0,1]$. ∎
Proposition 5.3.
@@
(5.7) $\begin{split}\frac{d}{dk}\left(\frac{\beta_{0}}{\beta_{1}}\right)&=-\frac{(d_{1}-1)(d_{2}-2)}{\beta_{1}^{2}}\tilde{\beta},\end{split}$
@@
(5.8) $\begin{split}\tilde{\beta}&=(2d_{1}^{2}d_{2}^{2}+2d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}+d_{2}^{2})k^{2}+(4d_{1}^{3}d_{2}+4d_{1}^{2}d_{2}^{2}-6d_{1}^{2}d_{2}-4d_{1}d_{2}^{2}+2d_{1}d_{2})k\\ &\quad+2d_{1}^{4}+2d_{1}^{3}d_{2}-4d_{1}^{3}-3d_{1}^{2}d_{2}+2d_{1}^{2}+d_{1}d_{2}.\end{split}$
The factor $\tilde{\beta}$ is positive as each coefficient for $k^{i}$ is positive given $d_{2}\geq d_{1}\geq 2$. Hence, we have $\min\limits_{k\in[0,1]}\left\{\frac{\beta_{0}}{\beta_{1}}\right\}=\frac{\beta_{0}}{\beta_{1}}(1)=\frac{1}{n+d_{1}}$. Note that if $(d_{1},d_{2})=(2,2)$, we have $\beta_{0}(k)=-2k^{2}+8k+6=\frac{1}{6}\beta_{1}(k)$. ∎
Proposition 5.4.
For any $d_{2}\geq d_{1}\geq 2$, we have $\beta_{2}(k)>0$ for any $k\in[0,1]$.
@@
(5.9) $\begin{split}\beta_{2}(k)&=(d_{1}-1)(d_{2}-1)d_{2}^{2}k^{3}+d_{1}d_{2}^{2}(k^{2}-k^{3})+(2d_{1}^{2}d_{2}^{2}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+2d_{1}d_{2})k^{2}\\ &\quad+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k+2d_{1}^{3}(1-k)\\ &>(2d_{1}^{2}d_{2}^{2}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+2d_{1}d_{2})k^{2}+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k\\ &\geq(-4d_{1}^{2}d_{2}+2d_{1}d_{2})k^{2}+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k\\ &\geq(d_{1}^{3}d_{2}+2d_{1}d_{2}-3d_{1}^{2}d_{2})k\\ &\geq 0.\end{split}$
@@
(5.10) $\theta_{2}=4d_{1}^{4}(d_{1}+1)^{2}\theta_{2a}\theta_{2b},\quad\theta_{1}=\theta_{1a}\theta_{1b},\quad\theta_{0}=\theta_{0a}\theta_{0b},$
@@
(5.11) $\begin{split}\theta_{2a}(k)&=(d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-4d_{1}^{2}+4d_{1}d_{2})k+d_{1}^{3}d_{2}-d_{1}^{2}d_{2}+4d_{1}^{2}\\ \theta_{2b}(k)&=(2d_{1}^{2}d_{2}^{2}-d_{1}d_{2}^{3}+d_{2}^{3}-d_{2}^{2})k^{3}\\ &\quad+(4d_{1}^{3}d_{2}-4d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}+4d_{1}d_{2}^{2}-2d_{1}d_{2})k^{2}\\ &\quad+(2d_{1}^{4}-5d_{1}^{3}d_{2}-2d_{1}^{3}+5d_{1}^{2}d_{2})k-2d_{1}^{4}+2d_{1}^{3}\\ \theta_{1a}(k)&=-4d_{1}^{2}d_{2}(d_{2}-1)^{2}(d_{1}+1)\\ \theta_{1b}(k)&=(d_{1}^{3}d_{2}^{5}-d_{1}^{2}d_{2}^{5}-4d_{1}^{3}d_{2}^{3}+3d_{1}^{2}d_{2}^{4}-d_{1}d_{2}^{4}+2d_{1}d_{2}^{3}-d_{2}^{4})k^{5}\\ &\quad+(4d_{1}^{4}d_{2}^{4}+4d_{1}^{4}d_{2}^{3}-6d_{1}^{3}d_{2}^{4}-16d_{1}^{4}d_{2}^{2}+18d_{1}^{3}d_{2}^{3}+2d_{1}^{2}d_{2}^{4}\\ &\quad\quad+4d_{1}^{3}d_{2}^{2}-10d_{1}^{2}d_{2}^{3}+8d_{1}^{2}d_{2}^{2}-6d_{1}d_{2}^{3})k^{4}\\ &\quad+(6d_{1}^{5}d_{2}^{3}+12d_{1}^{5}d_{2}^{2}-14d_{1}^{4}d_{2}^{3}-20d_{1}^{5}d_{2}+35d_{1}^{4}d_{2}^{2}+7d_{1}^{3}d_{2}^{3}\\ &\quad\quad+12d_{1}^{4}d_{2}-29d_{1}^{3}d_{2}^{2}+d_{1}^{2}d_{2}^{3}+8d_{1}^{3}d_{2}-12d_{1}^{2}d_{2}^{2})k^{3}\\ &\quad+(4d_{1}^{6}d_{2}^{2}+12d_{1}^{6}d_{2}-16d_{1}^{5}d_{2}^{2}-8d_{1}^{6}+28d_{1}^{5}d_{2}+8d_{1}^{4}d_{2}^{2}\\ &\quad\quad+8d_{1}^{5}-32d_{1}^{4}d_{2}+4d_{1}^{3}d_{2}^{2}-8d_{1}^{3}d_{2})k^{2}\\ &\quad+(d_{1}^{7}d_{2}+4d_{1}^{7}-9d_{1}^{6}d_{2}+8d_{1}^{6}+3d_{1}^{5}d_{2}-12d_{1}^{5}+5d_{1}^{4}d_{2})k\\ &\quad-2d_{1}^{7}+2d_{1}^{5}\\ \theta_{0a}(k)&=kd_{2}^{2}(d_{2}-1)^{4}(2d_{1}+d_{2}k)((2d_{1}d_{2}-2d_{1}+d_{2})k+2d_{1}^{2}+2d_{1})\\ \theta_{0b}(k)&=(2d_{1}^{2}-1)d_{2}^{2}k^{2}+(4d_{1}^{3}-2d_{1}^{2}-2d_{1})d_{2}k+2d_{1}^{3}(d_{1}-1).\end{split}$
@@
We use Maple to obtain the formula. ∎
Proposition 5.6.
For any $d_{2}\geq d_{1}\geq 2$, we have $\theta_{2}<0$ and $\theta_{0}\geq 0$ on $[0,1]$. The function $\theta_{0}$ only vanishes at $k=0$.
@@
(5.12) $\begin{split}\theta_{2a}&=d_{2}^{2}(d_{1}d_{2}-d_{1}-d_{2})k^{2}+(d_{2}^{2}-d_{1}d_{2})k^{2}+(d_{1}d_{2}^{2}+d_{1}d_{2})(k-k^{2})+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2})k\\ &\quad+(2d_{1}^{2}d_{2}-d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+3d_{1}d_{2}-3d_{1}^{2})k+(d_{1}^{2}d_{2}+d_{1}^{2})(1-k)+(d_{1}^{3}d_{2}-2d_{1}^{2}d_{2}+3d_{1}^{2}).\end{split}$
Given $d_{2}\geq d_{1}\geq 2$, each term above is non-negative on $[0,1]$. In particular, the last term is positive. Hence $\theta_{2a}>0$ on $[0,1]$.
@@
(5.13) $\begin{split}\theta_{2b}&=-(d_{1}d_{2}^{3}-d_{2}^{3}+d_{2}^{2})k^{3}-2d_{1}^{2}d_{2}^{2}(k^{2}-k^{3})\\ &\quad-(d_{1}^{3}d_{2}-3d_{1}^{2}d_{2}+2d_{1}d_{2})k^{2}-(2d_{1}^{2}d_{2}^{2}-4d_{1}d_{2}^{2})k^{2}\\ &\quad-(5d_{1}^{3}d_{2}-5d_{1}^{2}d_{2})(k-k^{2})-(2d_{1}^{4}-2d_{1}^{3})(1-k)\\ \end{split}.$
Given $d_{2}\geq d_{1}\geq 2$, all terms above are non-positive and do not vanish simultaneously for any $k\in[0,1]$. We have $\theta_{2b}<0$ on $[0,1]$. The proof is complete. ∎
Acknowledgement. The author would like to thank the conference organizers of “Einstein Spaces and Special Geometry” at the Institut Mittag-Leffler in July 2023, where part of this work was presented. The author thanks Christoph Böhm, Andrew Dancer, Lorenzo Foscolo, and Claude LeBrun for insightful discussions during the conference. Special thanks to Christoph Böhm for his advice in proving the inhomogeneity. The author also thanks McKenzie Wang and Xiping Zhu for valuable comments on an early draft of this paper.

### v5

@@ Appendix B Computations of Rational functions
[…]
The last four terms above are non-negative and do not vanish simultaneously for any $k\in[0,1]$. The coefficient of $k^{3}$ is also non-negative since
$d_{1}^{2}d_{2}^{2}+d_{1}d_{2}^{3}-2d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-d_{2}^{3}+d_{1}d_{2}+d_{2}^{2}=(d_{2}-2)(d_{1}d_{2}(d_{1}+d_{2})-d_{2}(d_{2}+1))+(d_{1}-2)d_{2}\geq 0.$
Therefore, given $d_{1},d_{2}\geq 2$, the polynomial above is non-negative on $[0,1]$. Hence, we have $P_{X}+Q_{X}<0$ on $[0,1]$. ∎
Proposition B.2.
For any $d_{1},d_{2}\geq 2$, the inequalities $\beta_{0}(k)>0$ and $\beta_{1}(k)>0$ hold for any $k\in[0,1]$.
@@
[…]
Since
(B.4) $\begin{split}\beta_{1}(0)&=2d_{1}^{4}+d_{1}^{3}d_{2}-2d_{1}^{2}d_{2}+2d_{1}^{2}+d_{1}d_{2}-4d_{1}\\ &\geq 2d_{1}^{4}+d_{1}d_{2}\\ &>0\end{split}$
and
(B.5) $\begin{split}\frac{d\beta_{1}}{dk}(0)&=4d_{1}^{3}d_{2}+2d_{1}^{2}d_{2}^{2}-4d_{1}^{3}-2d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+4d_{1}d_{2}+d_{2}^{2}+4d_{1}-2d_{2}\\ &\geq 4d_{1}^{3}d_{2}+2d_{1}^{2}d_{2}^{2}-4d_{1}^{3}-2d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+4d_{1}d_{2}+4d_{1}\\ &=(3d_{1}^{3}d_{2}-4d_{1}^{3})+(d_{1}^{3}d_{2}-2d_{1}^{2}d_{2})+(2d_{1}^{2}d_{2}^{2}-3d_{1}d_{2}^{2})+4d_{1}d_{2}+4d_{1}\\ &>0\end{split}$
if $d_{1},d_{2}\geq 2$, it is clear that $\beta_{1}>0$ on $[0,1]$ if the coefficient of $k^{2}$ is non-negative. Since
$\beta_{0}(1)=(d_{1}-1)(d_{1}+d_{2})(d_{1}+d_{2}-1)>0,$
we have $\beta_{1}>0$ on $[0,1]$ even if the coefficient of $k^{2}$ is negative. ∎
Proposition B.3.
@@
[…]
@@
(B.7) $\begin{split}\tilde{\beta}&=(2d_{1}^{2}d_{2}^{2}+2d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}+d_{2}^{2})k^{2}+(4d_{1}^{3}d_{2}+4d_{1}^{2}d_{2}^{2}-6d_{1}^{2}d_{2}-4d_{1}d_{2}^{2}+2d_{1}d_{2})k\\ &\quad+2d_{1}^{4}+2d_{1}^{3}d_{2}-4d_{1}^{3}-3d_{1}^{2}d_{2}+2d_{1}^{2}+d_{1}d_{2}.\end{split}$
The factor $\tilde{\beta}$ is positive as each coefficient for $k^{i}$ is positive given $d_{1},d_{2}\geq 2$. Hence, we have $\min\limits_{k\in[0,1]}\left\{\frac{\beta_{0}}{\beta_{1}}\right\}=\frac{\beta_{0}}{\beta_{1}}(1)=\frac{1}{n+d_{1}}$. Note that if $(d_{1},d_{2})=(2,2)$, we have $\beta_{0}(k)=-2k^{2}+8k+6=\frac{1}{6}\beta_{1}(k)$. ∎
Proposition B.4.
For any $d_{1},d_{2}\geq 2$, we have $\beta_{2}(k)>0$ for any $k\in[0,1]$.
@@
(B.8) $\begin{split}\beta_{2}(k)&=(d_{1}-1)(d_{2}-1)d_{2}^{2}k^{3}+d_{1}d_{2}^{2}(k^{2}-k^{3})+(2d_{1}^{2}d_{2}^{2}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+2d_{1}d_{2})k^{2}\\ &\quad+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k+2d_{1}^{3}(1-k)\\ &>(2d_{1}^{2}d_{2}^{2}-4d_{1}^{2}d_{2}-3d_{1}d_{2}^{2}+2d_{1}d_{2})k^{2}+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k\\ &\geq(-4d_{1}^{2}d_{2}+2d_{1}d_{2})k^{2}+(d_{1}^{3}d_{2}+d_{1}^{2}d_{2})k\\ &\geq(d_{1}^{3}d_{2}+2d_{1}d_{2}-3d_{1}^{2}d_{2})k\\ &\geq 0.\end{split}$
@@
(B.9) $\theta_{2}=4d_{1}^{4}(d_{1}+1)^{2}\theta_{2a}\theta_{2b},\quad\theta_{1}=\theta_{1a}\theta_{1b},\quad\theta_{0}=\theta_{0a}\theta_{0b},$
@@
(B.10) $\begin{split}\theta_{2a}(k)&=(d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-4d_{1}^{2}+4d_{1}d_{2})k+d_{1}^{3}d_{2}-d_{1}^{2}d_{2}+4d_{1}^{2},\\ \theta_{2b}(k)&=(2d_{1}^{2}d_{2}^{2}-d_{1}d_{2}^{3}+d_{2}^{3}-d_{2}^{2})k^{3}\\ &\quad+(4d_{1}^{3}d_{2}-4d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}+4d_{1}d_{2}^{2}-2d_{1}d_{2})k^{2}\\ &\quad+(2d_{1}^{4}-5d_{1}^{3}d_{2}-2d_{1}^{3}+5d_{1}^{2}d_{2})k-2d_{1}^{4}+2d_{1}^{3},\\ \theta_{1a}(k)&=-4d_{1}^{2}d_{2}(d_{2}-1)^{2}(d_{1}+1),\\ \theta_{1b}(k)&=(d_{1}^{3}d_{2}^{5}-d_{1}^{2}d_{2}^{5}-4d_{1}^{3}d_{2}^{3}+3d_{1}^{2}d_{2}^{4}-d_{1}d_{2}^{4}+2d_{1}d_{2}^{3}-d_{2}^{4})k^{5}\\ &\quad+(4d_{1}^{4}d_{2}^{4}+4d_{1}^{4}d_{2}^{3}-6d_{1}^{3}d_{2}^{4}-16d_{1}^{4}d_{2}^{2}+18d_{1}^{3}d_{2}^{3}+2d_{1}^{2}d_{2}^{4}\\ &\quad\quad+4d_{1}^{3}d_{2}^{2}-10d_{1}^{2}d_{2}^{3}+8d_{1}^{2}d_{2}^{2}-6d_{1}d_{2}^{3})k^{4}\\ &\quad+(6d_{1}^{5}d_{2}^{3}+12d_{1}^{5}d_{2}^{2}-14d_{1}^{4}d_{2}^{3}-20d_{1}^{5}d_{2}+35d_{1}^{4}d_{2}^{2}+7d_{1}^{3}d_{2}^{3}\\ &\quad\quad+12d_{1}^{4}d_{2}-29d_{1}^{3}d_{2}^{2}+d_{1}^{2}d_{2}^{3}+8d_{1}^{3}d_{2}-12d_{1}^{2}d_{2}^{2})k^{3}\\ &\quad+(4d_{1}^{6}d_{2}^{2}+12d_{1}^{6}d_{2}-16d_{1}^{5}d_{2}^{2}-8d_{1}^{6}+28d_{1}^{5}d_{2}+8d_{1}^{4}d_{2}^{2}\\ &\quad\quad+8d_{1}^{5}-32d_{1}^{4}d_{2}+4d_{1}^{3}d_{2}^{2}-8d_{1}^{3}d_{2})k^{2}\\ &\quad+(d_{1}^{7}d_{2}+4d_{1}^{7}-9d_{1}^{6}d_{2}+8d_{1}^{6}+3d_{1}^{5}d_{2}-12d_{1}^{5}+5d_{1}^{4}d_{2})k\\ &\quad-2d_{1}^{7}+2d_{1}^{5},\\ \theta_{0a}(k)&=kd_{2}^{2}(d_{2}-1)^{4}(2d_{1}+d_{2}k)((2d_{1}d_{2}-2d_{1}+d_{2})k+2d_{1}^{2}+2d_{1}),\\ \theta_{0b}(k)&=(2d_{1}^{2}-1)d_{2}^{2}k^{2}+(4d_{1}^{3}-2d_{1}^{2}-2d_{1})d_{2}k+2d_{1}^{3}(d_{1}-1).\end{split}$
@@
We use Maple to obtain the formulas. ∎
Proposition B.6.
For any $d_{1},d_{2}\geq 2$, we have $\theta_{2}<0$ and $\theta_{0}\geq 0$ on $[0,1]$. The function $\theta_{0}$ only vanishes at $k=0$.
@@
(B.11) $\begin{split}\theta_{2a}&=(d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}-4d_{1}^{2}+4d_{1}d_{2})k+d_{1}^{3}d_{2}-d_{1}^{2}d_{2}+4d_{1}^{2}\\ &=(d_{1}d_{2}^{3}-2d_{1}d_{2}^{2}-d_{2}^{3}-2d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}-2d_{1}d_{2}^{2}+4d_{1}d_{2})k+d_{1}^{3}d_{2}-d_{1}^{2}d_{2}+4d_{1}^{2}(1-k)\\ &\geq(d_{1}d_{2}^{3}-d_{2}^{3}-2d_{1}d_{2}+d_{2}^{2})k^{2}\\ &\quad+(2d_{1}^{2}d_{2}^{2}-2d_{1}^{2}d_{2}-4d_{1}d_{2}^{2}+4d_{1}d_{2})k+4d_{1}^{2}(1-k)+d_{1}^{2}d_{2}(d_{1}-1).\end{split}$
Given $d_{1},d_{2}\geq 2$, each term above is non-negative on $[0,1]$. In particular, the last term is positive. Hence $\theta_{2a}>0$ on $[0,1]$.
@@
(B.12) $\begin{split}\theta_{2b}&=-(d_{1}d_{2}^{3}-d_{2}^{3}+d_{2}^{2})k^{3}-2d_{1}^{2}d_{2}^{2}(k^{2}-k^{3})\\ &\quad-(d_{1}^{3}d_{2}-3d_{1}^{2}d_{2}+2d_{1}d_{2})k^{2}-(2d_{1}^{2}d_{2}^{2}-4d_{1}d_{2}^{2})k^{2}\\ &\quad-(5d_{1}^{3}d_{2}-5d_{1}^{2}d_{2})(k-k^{2})-(2d_{1}^{4}-2d_{1}^{3})(1-k).\end{split}$
Given $d_{1},d_{2}\geq 2$, all terms above are non-positive and do not vanish simultaneously for any $k\in[0,1]$. We have $\theta_{2b}<0$ on $[0,1]$. The proof is complete. ∎
Appendix C Examples
Recall in Section 4, we have $\chi_{d_{1},d_{2}}=\left\{\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)},A_{1},A_{2}\right\}$. The constant $A_{1}$ is the minimum of the rational function $F$ in Proposition 4.15, which can be explicitly computed with a fixed $(d_{1},d_{2})$. On the other hand, the constant $A_{2}$ is the only positive root of the quadratic function $\Theta_{d_{1},d_{2}}(A,k)$ as in Proposition B.5. Instead of computing $A_{2}$, it suffices to check $\Theta_{d_{1},d_{2}}(A,k)>0$ for any $k\in[0,1]$. Following the guideline described above, we present the following two examples.
Example C.1.
Consider the twistor bundle over $\mathbb{HP}^{1}$, whose group triple is
$(\mathsf{K},\mathsf{H},\mathsf{G})=(Sp(1)U(1)U(1),Sp(1)Sp(1),Sp(2)Sp(1)).$
The associated structural triple is $(2,4,1)$.
For $(d_{1},d_{2})=(2,4)$, we have
$F=\frac{9\left(16k^{3}+16k^{2}+32k+16\right)}{\left(4k+4\right)\left(48k+32\right)}.$
Since
$F^{\prime}=\frac{9\left(k^{2}+3k+1\right)\left(3k^{2}+k-1\right)}{4\left(k+1\right)^{2}\left(3k+2\right)^{2}}$
vanishes at $k_{*}=-\frac{1}{6}+\frac{\sqrt{13}}{6}\approx 0.434$ we have $A_{1}=\min\limits_{k\in[0,1]}F=F(k_{*})=\frac{19\sqrt{13}+47}{56+16\sqrt{13}}\approx 1.016>1$
On the other hand, we have
(C.1) $\begin{split}\Theta_{2,4}\left(1,k\right)&=1327104k^{5}+1658880k^{4}+663552k^{3}+221184k^{2}+147456k+36864>0.\end{split}$
Hence $A_{2}>1$. With $\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}=\frac{9}{8}>1$, we know that $1<\chi_{2,4}$. The Einstein metric on $F^{2}$ predicted in [GPP90, Böh98, DHW13] is recovered.
Example C.2.
Consider the quaternionic Hopf fibration over $\mathbb{HP}^{1}$, whose group triple is
$(\mathsf{K},\mathsf{H},\mathsf{G})=(Sp(1)\triangle Sp(1),Sp(1)Sp(1)Sp(1),Sp(2)Sp(1)).$
The associated structural triple is $\left(3,4,\frac{1}{4}\right)$.
For $(d_{1},d_{2})=(3,4)$, we have
$F=\frac{4\left(48k^{3}+72k^{2}+90k+54\right)}{\left(4k+6\right)\left(24k^{2}+132k+108\right)}.$
Since
$F^{\prime}=\frac{2\left(2k^{2}+6k+3\right)\left(44k^{2}+42k-9\right)}{\left(2k+3\right)^{2}\left(2k+9\right)^{2}\left(k+1\right)^{2}}$
vanishes at $k_{*}=-\frac{21}{44}+\frac{3\sqrt{93}}{44}\approx 0.180$ we have $A_{1}=\min\limits_{k\in[0,1]}F_{1}=F_{1}(k_{*})=\frac{503\sqrt{93}+2331}{10785+1159\sqrt{93}}\approx 0.327>\frac{1}{4}$
On the other hand, we have
(C.2) $\begin{split}\Theta_{3,4}\left(\frac{1}{4},k\right)&=9082368k^{5}+29673216k^{4}+38351232k^{3}+24774336k^{2}+8398080k+1259712>0.\end{split}$
Hence $A_{2}>\frac{1}{4}$. With $\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}=\frac{1}{3}$, we know that $\frac{1}{4}<\chi_{3,4}$. The Einstein metric on $\mathbb{HP}^{2}\sharp\overline{\mathbb{HP}}^{2}$ in [Böh98] is recovered.
Acknowledgement. The author would like to thank the organizers of the conference “Einstein Spaces and Special Geometry” at the Institut Mittag-Leffler in July 2023, where part of this work was presented. The author thanks Christoph Böhm, Andrew Dancer, Lorenzo Foscolo, and Claude LeBrun for insightful discussions during the conference. Special thanks to Christoph Böhm for helpful suggestions that improved the exposition and overall structure of the paper. The author also thanks McKenzie Wang and Xiping Zhu for valuable comments on an early draft of this paper.